# Multi-Robot Navigation in Social Mini-Games: Definitions, Taxonomy, and Algorithms

### v3

@@ 3.2.2 Communication
Communication among agents plays a key role in understanding the path planner for deadlock avoidance. Agents can either be Communicated or Uncommunicated, depending on whether they share information with each other during decision-making. When agents are communicated, they share information and coordinate their actions. By sharing their plans, intentions, and observations, Communicated agents can work together more effectively. This helps in avoiding conflicts and ensuring smoother navigation, as all agents are aware of each other’s actions and can adjust accordingly. The centralized coordinated methods are generally communicated because the agents have to communicate with the central coordinator/agent for decision-making. In contrast, Uncommunicated agents do not share information with each other. Each agent makes decisions independently, based only on its own observations and goals. While this approach simplifies the communication requirements, it can lead to more conflicts and less coordinated movements, and raise potential deadlocks among agents. However, with a well-designed control strategy, uncommunicated agents can still navigate effectively, relying on their ability to adapt to changing conditions and other agents’ actions. The decentralized methods generally behave in an uncommunicated manner, where each agent makes its own independent decision to generate the controls and methods like IMPC-DR [77] and ORCA-MAPF [46] come under this category. Fig 7 and 8 categorize the various papers under Communicated and Uncommunicated, respectively.
@@ 3.2.3 Deadlock Handling
In multi-robot social navigation, how an agent deals with deadlocks is very important. Deadlock often refers to situations where robot states converge to an undesirable equilibrium that causes the robots to stall [103, 128, 129]. Deadlock handling can be divided into two main types: Prevention and Resolution. Generally, there are two main approaches: Proactive (Deadlock Prevention) and Reactive (Deadlock Resolution), each with its own style and advantages. The Proactive Strategy involves planning ahead. Agents using this approach constantly analyze their environment to predict and prevent potential deadlocks before they happen. This is like a chess player who thinks several moves ahead. Examples of this approach include IMPC-DR [77] and MERRY-GO-Round [130] that combine deadlock resolution strategies with trajectory planning and navigation to ensure smoother movements, thus avoiding potential deadlocks. However, this requires a good understanding and prediction of the environment, as wrong predictions can lead to unnecessary detours or inefficient paths. The Reactive Strategy, on the other hand , focuses on immediate response. Agents using this method stay alert for potential deadlocks as they move. They don’t act or react until a deadlock is about to happen. Once they detect it, the robots quickly respond to resolve or avoid the deadlock by taking temporarily adjusted trajectories. This strategy deals with problems as they come up. Examples of this approach are methods that include the right-hand rule [64], auctions [131], ORCA-MAPF [46], and adaptive rotational strategies [128, 129]. The Reactive Strategy is flexible and can adapt to changes, but it might struggle if many problems occur at once. Fig. 9 and 10 categorize the MRNs for Deadlock Resolution and Prevention, respectively.
@@ 3.2.4 Invasiveness
In multi-robot social navigation, how much an agent affects its environment is crucial. Agents can be either Invasive or Non-Invasive. Invasive agents significantly change their environment or the paths of other agents to avoid deadlocks. They may alter their own path drastically or influence the movement of other agents. This approach ensures that the agent can navigate smoothly, but it can disrupt the environment and other agents’ plans. IMPC-DR [77] and Right Hand Rule [64] are the primary example for this category where the agents makes the other agent change their path and change its own path in a more exhaustive way to prevent/resolve the deadlock. Non-invasive agents, on the other hand, try to navigate without making significant changes to their environment or affecting other agents. They stick to predefined paths and make minimal adjustments to avoid conflicts. This approach reduces disruption and is more predictable, but it might be less flexible in handling complex or dynamic environments. The foremost examples in this category are Auctions and ORCA-MAPF [46] where the agents adjusts their velocity to avoid potential deadlocks and collisions without deviating from their path. Fig. 11 and 12 categorizes the different MRN algorithms in the Invasive and Non-invasive categories, respectively.
@@ 3.2.5 Cooperation
In multi-robot social navigation, cooperation refers to how agents work together to achieve smooth and efficient movement. Cooperation can be characterized based on whether agents share the same/different objectives (cost functions) and whether they use the same or different planning strategies. In fully Cooperative settings, agents optimize a common cost function, working toward shared goals through coordinated planning. Methods like PRIMAL [132] and CBM [133] follow this style, using explicit communication or negotiation to achieve joint plans. In Non-Cooperative settings, agents may have distinct cost functions and prioritize their own objectives, requiring implicit cooperation through local observations. Implicit (Semi) cooperation methods, such as ORCA-MAPF and Reciprocal Velocity Obstacles (RVO) [134], adaptively adjust trajectories based on predicted collisions without explicit communication. This relationship between cost and planning strategies (algorithms) are organized as shown in Fig 13 and Fig 14. Understanding these structures is key to designing socially compliant and deadlock-free multi-robot systems.
@@ 3.2.6 Observability
In multi-robot social navigation, observability describes how much information an agent has about other agents in the environment. Different levels of observability influence how agents plan and coordinate their actions. Under Full Observability, agents have complete knowledge of all other agents’ states, including positions, velocities, and goals, enabling highly coordinated planning, as seen in centralized methods like CBS [44], ECBS [65] and other MAPF solvers (Categorization in Fig. 15). In Local Observability, agents rely only on nearby sensing to detect others, as used in decentralized methods such as ORCA, Reciprocal Velocity Obstacles (RVO) [47], and SIPP [149] (see Fig. 16). Predicted Observability refers to agents forecasting others’ future movements using internal models, often applied in motion prediction frameworks like Social-LSTM [150]. This is where the agents are fully aware of the environment, but partially aware of other agents. Belief-space observability involves maintaining probabilistic beliefs over others’ states or intents, which is commonly addressed through POMDP-based planners [151] or uncertainty-aware imitation learning. Finally, under Minimal Observability, agents do not explicitly model others and treat them as static or reactive obstacles, simplifying planning but limiting social awareness. In this case, the agents are neither aware of other agents nor the environment. Choosing the appropriate observability level is crucial for balancing system complexity, coordination ability, and robustness in dynamic social environments.
@@ 3.3.1 Safety
Provable safety can be achieved by single-integrator systems e.g. ORCA framework from Van Den Berg et al. [134] and its non-holonomic variant [48], which are effective for fast and exact multi-robot navigation. ORCA conservatively imposes collision avoidance constraints on the motion of a robot in terms of half-planes in the space of velocities. The optimal collision-free velocity can then be quickly found by solving a linear program. Proving safety is harder for systems with double-integrator dynamics, therefore safety in these systems depends on the planning frequency of the system. For example, the NH-TTC algorithm [153] uses gradient descent to minimize a cost function comprising a goal reaching term and a time-to-collision term, which rises to infinity as the agent approaches immediate collision. NH-TTC guarantees safety in the limit as the planning frequency approaches infinity. Other optimization-based approaches use Model Predictive Control(MPC); in such approaches, safety depends not only on the planning frequency but also on the length of the planning horizon. Finally, Control Barrier Functions (CBFs) [64] can be used to design controllers that can guarantee safety via the notion of forward invariance of a set i.e. if an agent starts out in a safe set at the initial time step, then it remains safe for all future time steps, that is, it will never leave the safe set. CBFs can also be extended to a probabilistic setting [154] to account for realistic uncertainties in system dynamics and observations.
@@ 3.3.2 Liveness
Liveness is a property that means robots are continuously making progress towards their goals, and only stop once they have reached their goals. In general MRN, liveness is typically considered a “soft constraint” as opposed to safety, which is often a hard constraint. However, MRN in SMGs often result in deadlocks, so they become critical. Methods for ensuring liveness are characterized along the taxonomy can either be deadlock-preventing [14, 155, 156] or deadlock resolving [64, 141, 103, 146] (previously discussed in Sec. 3.2.3). The difference lies in when they solve a deadlock. Deadlock prevention methods are generally smoother and less invasive as they do not wait for a deadlock to happen in the first place. Next, certain liveness methods rely on agents communicating between themselves to resolve conflicts [93, 157, 94]. Finally, some liveness methods use global planning methods to plan non-colliding paths [46].
3.3.3 Fairness
Fairness is an important property in SMGs that takes into account the priorities of the agents, unlike safety or liveness. Fairness aims to maximize the social welfare of the entire system. For example, consider an ambulance robot and a delivery robot arrive at an intersection. From a safety and liveness perspective, there are two solutions (ambulance goes first followed by the delivery robot, and vice-versa). However, the ambulance clearly has the right of way, and so from a fairness perspective, the only correct solution is for the ambulance robot to go through first. Fairness is a difficult property to achieve as the individual priority levels are private and are typically unknown to other robots, and must instead be inferred. Fairness can be achieved through mechanisms that balance priorities [131, 158, 85], or fixed, agreed-upon rule-based decision-making such as the right-hand-rule [159, 64] that ensures all agents have a reasonable opportunity to make progress toward their goals. Fairness has also been studied as a credit-assignment problem in MARL [160, 161, 162, 163, 164].
@@ 3.3.4 Social Compliance
Social compliance refers to the ability of agents to navigate in a manner that respects social norms, conventions, and expectations within shared environments. Many simulators [32, 165] and datasets [29] have been proposed to develop methods that ensure social compliance. Social robot navigation is a vast area of research and we refer the reader to surveys in this area for a comprehensive treatment [34]. Strategies to promote social compliance include modeling human behavior through imitation learning [166, 167, 168, 169], rule-based systems that encode social conventions [170], and reinforcement learning approaches that reward socially acceptable actions [37, 171]. Recently, research has also started to study hybrid methods that combine imitation learning and classical methods [172].
@@ 3.3.5 Scalability
Scalability refers to an MRN solver’s ability to handle a growing number of robots without a loss in performance. This property is especially important in fields like warehouses and logistics, where many robot agents need to work together efficiently. Solvers that analytically prevent or resolve deadlocks are often constrained to a limited number of robots [147, 14, 146, 103]. On the other hand, multi-robot pathfinding algorithms operate in discrete space and discrete time and many efficient and fast solvers have been developed that are able to handle over thousands of robots [173]. We evaluate scalability in the SMG regime with respect to the active coupling set size , not total . Many solvers scale to large when SMGs are rare/small, but performance hinges on how cost/latency grow with during trigger activations (doorways, crossings). Reporting outcomes (SWG, TTF, IS, DR) stratified by separates algorithmic scaling from scene density.
4 Evaluation in SMGs
We first introduce a few methods that are evaluated for the different scenarios detailed in Section 2. Then we evaluate these methods using a benchmark software introduced in this work, called Social Mini-Games Library (SMGLib).
4.1 Representative SMG Solvers
Navigating from the structured domain of discrete MAPF algorithms to the domain of continuous approaches represents a significant evolution in multi-agent systems, especially within the context of SMGs. While the discrete algorithms, such as CBS [44] or ECBS [65], function within the confines of predetermined, often grid-based, spaces, continuous algorithms embrace the fluidity and intricacies of unquantized environments. This transition is akin to moving from a chessboard, with fixed squares and regimented moves, to an open field, where movements are unrestricted and can be infinitely varied. In SMGs, where nuanced interactions and dynamic adaptations are crucial, continuous algorithms provide the necessary finesse and granularity. They offer the potential for more natural and realistic agent behaviors, mirroring the seamless interactions observed in real-world social dynamics. As we venture into this continuous domain, we’ll explore strategies that harness this granularity, offering solutions that are not just mathematically efficient, but also socially coherent and intuitive. Figure 17 shows these methods graphically with multiple agents in action. In the following sections, we discuss inner-workings, strengths/limitations of some of the techniques, some of which were previously mentioned in Sec. 3.
4.1.1 CADRL [152]:
@@
Collision Avoidance with Deep Reinforcement Learning (CADRL) is a method that combines the fast decision-making of reaction-based approaches with the smooth motion of trajectory-based planning. It uses reinforcement learning to shift heavy computations from real-time execution to an offline training phase. This feature allows the system to learn a collision avoidance policy in advance, making real-time navigation faster and more efficient. A key challenge CADRL addresses is handling a changing number of agents in dynamic environments. To manage this, it uses Long Short-Term Memory (LSTM) networks to encode the varying information about nearby agents into a fixed-size vector. This enables CADRL to make decisions based on any number of surrounding agents without needing a fixed input size. By using LSTM, CADRL not only adapts to dynamic multi-agent scenarios but also retains important past information, leading to better decision-making. Overall, CADRL provides a scalable and efficient solution for real-time collision avoidance in complex, crowded environments.
4.1.2 Right-Hand-Rule (RHS) [64, 147]:

### v4

@@ 3.2.2 Communication
Communication among agents plays a key role in understanding the path planner for deadlock avoidance. Agents can either be Communicated or Uncommunicated, depending on whether they share information with each other during decision-making. When agents are communicated, they share information and coordinate their actions. By sharing their plans, intentions, and observations, Communicated agents can work together more effectively. This helps in avoiding conflicts and ensuring smoother navigation, as all agents are aware of each other’s actions and can adjust accordingly. The centralized coordinated methods are generally communicated because the agents have to communicate with the central coordinator/agent for decision-making. In contrast, Uncommunicated agents do not share information with each other. Each agent makes decisions independently, based only on its own observations and goals. While this approach simplifies the communication requirements, it can lead to more conflicts and less coordinated movements, and raise potential deadlocks among agents. However, with a well-designed control strategy, uncommunicated agents can still navigate effectively, relying on their ability to adapt to changing conditions and other agents’ actions. The decentralized methods generally behave in an uncommunicated manner, where each agent makes its own independent decision to generate the controls and methods like IMPC-DR [chen2023multi] and ORCA-MAPF [dergachev2021distributed] come under this category. Fig 9 and 10 categorize the various papers under Communicated and Uncommunicated, respectively.
@@ 3.2.3 Deadlock Handling
In multi-robot social navigation, how an agent deals with deadlocks is very important. Deadlock often refers to situations where robot states converge to an undesirable equilibrium that causes the robots to stall [grover2023before, zhang2025adaptive, zhang2025deadlock]. Deadlock handling can be divided into two main types: Prevention and Resolution. Generally, there are two main approaches: Proactive (Deadlock Prevention) and Reactive (Deadlock Resolution), each with its own style and advantages. The Proactive Strategy involves planning ahead. Agents using this approach constantly analyze their environment to predict and prevent potential deadlocks before they happen. This is like a chess player who thinks several moves ahead. Examples of this approach include IMPC-DR [chen2023multi] and MERRY-GO-Round [lee2025merry] that combine deadlock resolution strategies with trajectory planning and navigation to ensure smoother movements, thus avoiding potential deadlocks. However, this requires a good understanding and prediction of the environment, as wrong predictions can lead to unnecessary detours or inefficient paths. The Reactive Strategy, on the other hand , focuses on immediate response. Agents using this method stay alert for potential deadlocks as they move. They don’t act or react until a deadlock is about to happen. Once they detect it, the robots quickly respond to resolve or avoid the deadlock by taking temporarily adjusted trajectories. This strategy deals with problems as they come up. Examples of this approach are methods that include the right-hand rule [wang2017safety], auctions [chandra2023socialmapf], ORCA-MAPF [dergachev2021distributed], and adaptive rotational strategies [zhang2025adaptive, zhang2025deadlock]. The Reactive Strategy is flexible and can adapt to changes, but it might struggle if many problems occur at once. Fig. 11 and 12 categorize the MRNs for Deadlock Resolution and Prevention, respectively.
@@ 3.2.4 Invasiveness
Invasiveness is a property of an SMG solver that refers to the degree of change to an agent’s optimal or preferred trajectory induced by a control input during an SMG. For example, in the doorway example, in order to avoid a deadlock or collision, the agents must react. On the one extreme, a controller may just have an agent come to a complete stop or “go off to the side and wait”, yielding to the other agent. This is an example of an invasive solver. And if the controller simply tweaks the speed of the agent, without requiring it to deviate from its current trajectory, then the solver is said to be minimally invasive. We formalize this as follows:
Any solver that is not minimally invasive is generally termed as invasive to varying degrees. A minimally invasive perturbation does not cause a robot to deviate from its preferred trajectory (condition ()) only allowing it to speed up or slow down (condition ()). Invasive SMG solvers significantly change the agent’s environment or their paths to avoid deadlocks or collisions. They may also inadvertently alter the paths or influence the movements of other agents. This approach ensures that the agent can navigate smoothly, but it can disrupt the environment and other agents’ plans.
IMPC-DR [chen2023multi] and Right Hand Rule [wang2017safety] are the primary example for this category where a fixed rule-based approach necessitates agents to follow a specific order, which may cause agents who are moving fast but may have lower priority to slow down suddenly to prevent/resolve the deadlock. Examples of minimally invasive solvers include Auctions and ORCA-MAPF [dergachev2021distributed] where the agents adjust their speeds by the smallest amount needed to avoid potential deadlocks and collisions without deviating from their path. Fig. 13 and 14 categorizes the different MRN algorithms in the Invasive and Non-invasive categories, respectively.
@@ 3.2.5 Cooperation
Cooperation refers to how agents work together to achieve smooth and efficient movement. Cooperation can be characterized based on whether agents share the same/different objectives (cost functions) and whether they use the same or different planning strategies. In fully Cooperative settings, agents optimize a common cost function, working toward shared goals through coordinated planning. Methods like PRIMAL [everett2019primal] and CBM [ma2016optimal] follow this style, using explicit communication or negotiation to achieve joint plans. In Non-Cooperative settings, agents may have distinct cost functions and prioritize their own objectives, requiring implicit cooperation through local observations. Implicit (Semi) cooperation methods, such as ORCA-MAPF and Reciprocal Velocity Obstacles (RVO) [van2011reciprocal], adaptively adjust trajectories based on predicted collisions without explicit communication. This relationship between cost and planning strategies (algorithms) are organized as shown in Fig 15 and Fig 16. Understanding these structures is key to designing socially compliant and deadlock-free multi-robot systems.
@@ 3.2.6 Observability
Observability describes how much information an agent has about other agents in the environment. Different levels of observability influence how agents plan and coordinate their actions. Under Full Observability, agents have complete knowledge of all other agents’ states, including positions, velocities, and goals, enabling highly coordinated planning, as seen in centralized methods like CBS [sharon2015conflict], ECBS [barer2014suboptimal] and other MAPF solvers (Categorization in Fig. 17). In Local Observability, agents rely only on nearby sensing to detect others, as used in decentralized methods such as ORCA, Reciprocal Velocity Obstacles (RVO) [berg2008reciprocal], and SIPP [phillips2011sipp] (see Fig. 18). Predicted Observability refers to agents forecasting others’ future movements using internal models, often applied in motion prediction frameworks like Social-LSTM [alahi2016social]. This is where the agents are fully aware of the environment, but partially aware of other agents. Belief-space observability involves maintaining probabilistic beliefs over others’ states or intents, which is commonly addressed through POMDP-based planners [somani2013despot] or uncertainty-aware imitation learning. Finally, under Minimal Observability, agents do not explicitly model others and treat them as static or reactive obstacles, simplifying planning but limiting social awareness. In this case, the agents are neither aware of other agents nor the environment. Choosing the appropriate observability level is crucial for balancing system complexity, coordination ability, and robustness in dynamic social environments.
@@ 3.3.1 Safety
Provable safety can be achieved by single-integrator systems e.g. ORCA framework from Van Den Berg et al. [van2011reciprocal] and its non-holonomic variant [nh-orca], which are effective for fast and exact multi-robot navigation. ORCA conservatively imposes collision avoidance constraints on the motion of a robot in terms of half-planes in the space of velocities. The optimal collision-free velocity can then be quickly found by solving a linear program. Proving safety is harder for systems with double-integrator dynamics, therefore safety in these systems depends on the planning frequency of the system. For example, the NH-TTC algorithm [davis2019nh] uses gradient descent to minimize a cost function comprising a goal reaching term and a time-to-collision term, which rises to infinity as the agent approaches immediate collision. NH-TTC guarantees safety in the limit as the planning frequency approaches infinity. Other optimization-based approaches use Model Predictive Control(MPC); in such approaches, safety depends not only on the planning frequency but also on the length of the planning horizon. Finally, Control Barrier Functions (CBFs) [wang2017safety] can be used to design controllers that can guarantee safety via the notion of forward invariance of a set i.e. if an agent starts out in a safe set at the initial time step, then it remains safe for all future time steps, that is, it will never leave the safe set. CBFs can also be extended to a probabilistic setting [luo2020multi] to account for realistic uncertainties in system dynamics and observations.
@@ 3.3.2 Liveness
Liveness is a property that means robots are continuously making progress towards their goals, and only stop once they have reached their goals. In general MRN, liveness is typically considered a “soft constraint” as opposed to safety, which is often a hard constraint. However, MRN in SMGs often result in deadlocks, so they become critical. Methods for ensuring liveness are characterized along the taxonomy can either be deadlock-preventing [chandra2025deadlock, gouru2024livenet, chen2025livepoint] or deadlock resolving [wang2017safety, chen2024deadlock, grover2023before, grover2022semantically] (previously discussed in Sec. 3.2.3). The difference lies in when they solve a deadlock. Deadlock prevention methods are generally smoother and less invasive as they do not wait for a deadlock to happen in the first place. Next, certain liveness methods rely on agents communicating between themselves to resolve conflicts [garg2024deadlock, mahadevan2025gamechat, garg2024large]. Finally, some liveness methods use global planning methods to plan non-colliding paths [dergachev2021distributed].
3.3.3 Welfare Maximizing
Social welfare is an important property in SMGs that takes into account the priorities of the agents, along with the objective of the global system designer or society as a whole. Typically, social welfare is defined as a weighted linear combination of agents’ individual priorities as well as their local objective functions. For example, consider an ambulance robot and a delivery robot arrive at an intersection. From a safety and liveness perspective, there are two solutions (ambulance goes first followed by the delivery robot, and vice-versa). However, the ambulance clearly has the right of way, and so from a social welfare perspective, the only correct solution is for the ambulance robot to go through first. Welfare maximization is a difficult property to achieve as the individual priority levels are private and are typically unknown to other robots, and must instead be inferred. Additionally, maximizing welfare and individual agent priorities typically are at odds with each other. Welfare maximization can be achieved through mechanisms that balance priorities [chandra2023socialmapf, suriyarachchi2022gameopt, chandra2022gameplan], or fixed, agreed-upon rule-based decision-making such as the right-hand-rule [zhou2017fast, wang2017safety] that ensures all agents have a reasonable opportunity to make progress toward their goals. Welfare maximization has also been studied as a credit-assignment problem in MARL [nagpal2025leveraging, zhou2020learning, rahmattalabi2016d++, nguyen2018credit, feng2022multi].
@@ 3.3.4 Social Compliance
Social compliance refers to the ability of agents to navigate in a manner that respects social norms, conventions, and expectations within shared environments. Many simulators [sprague2023socialgym, holtz2022socialgym] and datasets [scanddataset] have been proposed to develop methods that ensure social compliance. Social robot navigation is a vast area of research and we refer the reader to surveys in this area for a comprehensive treatment [francis2025principles]. Strategies to promote social compliance include modeling human behavior through imitation learning [karnan2022voila, xiao2022learning, park2023learning, rosmann2017online], rule-based systems that encode social conventions [holtz2021iterative], and reinforcement learning approaches that reward socially acceptable actions [chen2017socially, chen2017decentralized]. Recently, research has also started to study hybrid methods that combine imitation learning and classical methods [raj2024rethinking].
@@ 3.3.5 Scalability
Scalability refers to an MRN solver’s ability to handle a growing number of robots without a loss in performance. This property is especially important in fields like warehouses and logistics, where many robot agents need to work together efficiently. Solvers that analytically prevent or resolve deadlocks are often constrained to a limited number of robots [wang2016safety, chandra2025deadlock, grover2022semantically, grover2023before]. On the other hand, multi-robot pathfinding algorithms operate in discrete space and discrete time and many efficient and fast solvers have been developed that are able to handle over thousands of robots [yan2025advancing]. We evaluate scalability in the SMG regime with respect to the active coupling set size , not total . Many solvers scale to large when SMGs are rare/small, but performance hinges on how cost/latency grow with during trigger activations (doorways, crossings). Reporting outcomes (SWG, TTF, IS, DR) stratified by separates algorithmic scaling from scene density.
4 Representative SMG Solvers
In this Section, we focus on a few specific approaches that have been been widely tested an adopted as baselines when solving SMGs. We acknowledge that this is an ongoing and emerging research area and this list is by no means exhaustive, but only serves to list the prominent approaches to serve as a starting point for new researches to the field. Figure 19 shows these methods graphically with multiple agents in action. In the following sections, we discuss inner-workings, strengths/limitations of some of the techniques, some of which were previously mentioned in Sec. 3.
4.1 CADRL [everett2021collision]:
@@
Collision Avoidance with Deep Reinforcement Learning (CADRL) is a method that combines the fast decision-making of reaction-based approaches with the smooth motion of trajectory-based planning. It uses reinforcement learning to shift heavy computations from real-time execution to an offline training phase. This feature allows the system to learn a collision avoidance policy in advance, making real-time navigation faster and more efficient. Each agent forms an observation from its own pose/velocity + goal and a set of nearby agents’ relative positions, velocities, and sizes/radii. The network outputs a real-time velocity command (often published as a commanded twist / linear+angular speed or planar velocity). A key challenge CADRL addresses is handling a changing number of agents in dynamic environments. To manage this, it uses Long Short-Term Memory (LSTM) networks to encode the varying information about nearby agents into a fixed-size vector. This enables CADRL to make decisions based on any number of surrounding agents without needing a fixed input size. By using LSTM, CADRL not only adapts to dynamic multi-agent scenarios but also retains important past information, leading to better decision-making. However, CADRL-style policies are learned and typically do not provide formal collision-free guarantees. Overall, CADRL provides a scalable and efficient solution for real-time collision avoidance in complex, crowded environments.
4.2 Right-Hand-Rule (RHS) [wang2017safety, wang2016safety]: